OpenStudy (anonymous):

find the quotients: 1: 4/7 divided by 7/8 = 2. 2 1/2 divided by 4=

OpenStudy (campbell_st):

the rule for dividing by a fraction is to flip the divisor and muliply so \[\frac{4}{7} \div \frac{7}{8} = \frac{4}{7} \times \frac{8}{7}\] now just evaluate in the the 2nd question the divisor of 4 can be written as \[4 = \frac{4}{1} \] so flip it and then multiply

OpenStudy (campbell_st):

if you have a mised number, 2 1/2 then make it an improper fraction